# A Tale of Two Imperatives: Privacy and Explainability

### v3

@@ 10 Related Work
Privacy-preserving machine learning (PPML). PPML is a critical component for RTP and includes a variety of techniques designed to safeguard individuals’ data while enabling meaningful machine learning analyses. Key methods in PPML include Differential Privacy (DP) Dwork et al. (2014); Abadi et al. (2016), Federated Learning (FL) McMahan et al. (2017); Kaissis et al. (2020); Konečný et al. (2016), Homomorphic Encryption (HE) Brand and Pradel (2023), and Secure Multi-Party Computation (SMPC) Zhou et al. (2024) to mention a few. DP is especially notable for ensuring that the inclusion or exclusion of any single data point has minimal impact on the results of computations, providing strong guarantees that sensitive information cannot be easily inferred from model outputs. FL enables decentralized training, allowing models to be trained on local devices without the need to share raw data. Homomorphic Encryption allows computations to be performed on encrypted data, while SMPC enables multiple parties to collaboratively train models without revealing their private data.
Among these approaches, DP is often regarded as the most suitable for many privacy-preserving applications due to its solid theoretical foundation and scalability. DP offers formal, quantifiable privacy guarantees, making it applicable across diverse domains. Unlike other methods, DP facilitates flexible trade-offs between privacy and utility by adjusting parameters like the noise scale, and it can be seamlessly integrated into a variety of machine learning algorithms. In comparison, techniques such as FL, HE, and SMPC offer privacy advantages but may introduce higher computational overheads and more complex infrastructure requirements Xu et al. (2021). These limitations can hinder their applicability in resource-constrained settings or in large-scale, real-time applications. Consequently, DP remains the preferred approach for balancing privacy protection with the efficient and scalable deployment of machine learning models. For a comprehensive review of PPML, we direct readers to the excellent survey by Boulemtafes et al. Boulemtafes et al. (2020).
Explainable machine learning (XAI). XAI is crucial for RTE. XAI methods are generally categorized into three types: local Ribeiro et al. (2016); Lundberg and Lee (2017); Sundararajan et al. (2017); Smilkov et al. (2017), global Ibrahim et al. (2019); Lundberg et al. (2020), and inherently interpretable models Ditz (2024); Molnar et al. (2020). Local explanations focus on providing insights into specific predictions for individual instances, while global explanations aim to clarify the overall behavior of the model across a range of examples. Inherently interpretable models, such as decision trees or linear models, are designed to be understandable from the outset. In this paper, we focus on local, post-hoc explainers (see Appendix C), which are particularly practical in industrial settings due to their “plug-and-play” nature. These methods do not require modifications to the model or its training process, making them highly adaptable in real-world applications Bhatt et al. (2020b).
These local explainers can be further divided into two categories: perturbation-based methods and gradient-based methods Saeed and Omlin (2023). Perturbation-based methods, such as LIME Ribeiro et al. (2016) and SHAP Lundberg and Lee (2017), analyze model predictions by altering the input and observing how the output changes. In contrast, gradient-based methods rely on the gradients of the model’s output with respect to the given inputs to generate explanations Saeed and Omlin (2023).
When evaluating the quality of explanations, key factors such as faithfulness, plausibility, and robustness are critical Zhou et al. (2021). Faithfulness refers to how accurately the explanation reflects the model’s actual decision-making process Lyu et al. (2024), while plausibility concerns how well the explanation aligns with human intuition or understanding Zhou et al. (2021); Lyu et al. (2024). Robustness measures the consistency of the explanation across small perturbations in the input or model Alvarez-Melis and Jaakkola (2018). Among these, faithfulness is widely regarded as the most important metric. An explanation that is not faithful to the model’s true behavior can mislead users, eroding trust in the system Lyu et al. (2024). While plausibility is important, it is insufficient by itself—an explanation that seems plausible to a human may still be unfaithful to the model’s actual behavior, which can ultimately deceive the end user Zhong et al. (2019). Therefore, faithfulness must remain the primary objective in XAI. However, measuring faithfulness presents its own challenges, and existing metrics are far from perfect, as discussed in Section 3.2. Even in domains like natural language processing, traditional erasure-based measures for evaluating faithfulness have proven flawed, leading to misleading conclusions about the quality of explanations, as noted in the recent work of Manna and Sett Manna and Sett (2024).
In this paper, we propose an alternative method for assessing the extent to which DP explanations can serve as a proxy for their non-DP counterparts. However, our mechanistic interpretation of DP models reveals that standard explainers are not well-suited for this context, due to the inherent differences between the nature of DP models and their non-DP counterparts.
However in this work, we did not consider model-agnostic perturbation-based explainers such as KernelSHAP, LIME, or Occlusion: other than their limited adaptation in computer vision tasks, Shokri et al. Shokri et al. (2021) criticized these methods for failing to follow the distribution of the training data and violating the data manifold hypothesis, which suggests that “data tend to lie near a low-dimensional manifold” Fefferman et al. (2016). Shokri et al. Shokri et al. (2021) further conjectured that this inherent shortcoming makes such methods fundamentally inappropriate for evaluating models with DP models. Subsequently, Kumar et al. Kumar et al. (2020) and Sundararajan et al. Sundararajan and Najmi (2020) also found that perturbation-based methods often operate outside the data’s true distribution, potentially leading to inaccurate or misleading explanations. While we did consider Grad-Shap, which can be seen as an approximation of Integrated Gradients by computing expectations of gradients over different baselines  [^21] Captum , we found that none of the selected explainers were effective or reliable when applied to DP models. This underscores the challenges of using traditional explainability techniques in the context of private models.
Privacy-preserving explainable AI (PPXAI). PPXAI methods have only recently begun to emerge, with approaches such as differentially private Locally Linear Maps (LLM) Harder et al. (2020), differentially private feature-based model explanations Patel et al. (2022), differentially private counterfactual explanations Mochaourab et al. (2021). For a comprehensive overview of PPXAI, we direct readers to the comprehensive survey by Nguyen et al. Nguyen et al. (2024). However, these methods are not included in our analysis due to their limited adoption and lack of extensive evaluation on critical aspects other than privacy preservation such as explanation quality and faithfulness, unlike the more widely used explainability techniques considered in this study.
@@
Recently, researchers have begun investigating the quality of explanations in privacy-preserving environments. Bozorgpanah et al. Bozorgpanah et al. (2022) generated private datasets (benchmark tabular datasets) using masking methods and compared the Shapley values Lundberg and Lee (2017) for test instances given the two models: trained on non private data and trained on private data. They found minimal effect on moderate protection. Lucieri et al. Lucieri et al. (2023) investigated the effect of DP training on concept-based explanations on biomedical datasets, and found that DP decreases average Concept Activation Vectors (CAV) accuracy and increases standard deviation; whereas Saifullah et al. (2024) has reported that differential privacy and federated learning may yield ‘noisy’ feature attribution score for post-hoc explainers. Berning et al. Berning et al. (2024) found that k-anonymity degrades the quality of counterfactual explanations on a tabular dataset. Bozorgpanah et al. Bozorgpanah and Torra (2024) generated private datasets (benchmark tabular datasets) using masking methods and noise addition and applied TreeSHAP to achieve plausible explanations.
Our work is distinct from all of these previous efforts, as we first establish the desiderata for DP explanations and then mechanistically explore why DP models are particularly challenging to explain using traditional post-hoc explainers. We not only investigate global DP and its inherent limitations for RTE but also explored local DP to generate proxy private explanations. While Lucieri et al. (2023) and Saifullah et al. (2024) focused on global DP, the rest of the existing literature has primarily worked with local DP and to the best of our knowledge we are the first one to mechanistically interpret the DP models and discovered why post-hoc explaienrs are inappropriate in this context.
@@
Analysing neural network similarity. Analysing neural network similarity is crucial for interpreting and improving model behavior. This can be broadly classified into two main types Klabunde et al. (2024): (i) representational similarity (RS), which measures differences in intermediate layer activations Kornblith et al. (2019); Cui et al. (2022); Raghu et al. (2017); Li et al. (2015), and (ii) functional similarity, which evaluates discrepancies in model outputs Hsu and Calmon (2022); Milani Fard et al. (2016); Marx et al. (2020).
@@
To investigate why DP models fail to accommodate common post-hoc explainers, we worked with RS to examine layer-wise changes in activations and their sensitivities. However, in such cases the similarity measures researchers employ adhere to specific invariances: Permutations, Orthogonal Transformations, Isotropic Scaling, Translations etc are to name a few. We have considered the mostly used set of invariances and chose CKA for measuring RS but CKA, by its design, comes with a few shortcomings as discussed in section 7. Consequently, we chose dCKA proposed by Cui et al. (2022) for our analysis. For a broader view on neural network similarity, we refer the readers to the wonderful survey by Klabunde et al. Klabunde et al. (2024).
11 Limitation of the Work
In this work, we primarily restricted our study to DP models, their activation patterns and their intrinsic behaviour. We have shown that an immediate consequence of the same is the non-functionality of famous gradient-based post-hoc explainers with the DP model. There can be different notions and setups for privacy-preserving machine learning and other classes of explainers are also there - we didn’t consider them for this research. Also, the (un)fairness of DP models Fioretto et al. (2022) is not considered under the scope of this research.
@@ 12 Conclusion and Future Studies
We started our research with a simple question of whether we can achieve RTP and RTE together. We kept an on eye the pitfalls of commonly used evaluation metrics for explainability and proposed our desiderata and found that no commonly used gradient-based explainers are useful for private models. We investigate the activations inside a DP model and how the model is sensitive towards those; we discovered that the intrinsic behaviour of DP models is the key reason behind this. Our ‘mechanistic’ Saphra and Wiegreffe (2024) insights of the private models across different privacy guarantees highlight how DP training alters the internal representations and their sensitivities. It gives a fresh perspective on interpreting DP models and their behaviour from a nuanced angle. Lastly, we make use of LDP to achieve private explanations and conclude our study by outlining the pipeline for the industrial software for our use case that respects both RTP and RTE.
In future studies, we aim to go deeper into the sensitivity landscape of DP models by investigating second-order and higher-order derivatives of the model’s output w.r.t. representations. While our current analysis focused on first-order gradients, examining higher-order derivatives could potentially uncover richer structural patterns in the sensitivity space that aren’t immediately apparent through first-order analysis. This exploration is particularly intriguing as higher-order derivatives might capture more nuanced interactions between features and reveal how privacy constraints affect these complex relationships. Specifically, we want to analyze how the Hessian and higher-order tensors of DP models differ from their non-private counterparts, which could provide insights into more complicated behavioural patterns. Understanding these higher-order sensitivities could not only enhance our mechanistic interpretation of DP models but also potentially lead to more sophisticated XAI method(s) that capture intricate model behaviours subject to privacy guarantees. This direction is especially promising as it might reveal whether the privacy-induced changes in model behaviour have deeper structural implications beyond what is observable through first-order sensitivity analysis.
@@ Acknowledgement
We are thankful to Jim Conant for the fruitful discussion on high-dimensional geometry. We are thankful to Yogesh Kumar, the co-author of dCKA Cui et al. (2022) for a productive discussion. We are also thankful to Dr. Venkata Abhinay Talasila from SRM University AP for his immense cooperation throughout.
Appendix A
Case: Normal, Grad-shap
Case: Normal, Integrated Gradients
Case: Normal, SmoothGrad
Case: Normal, Saliency
Case: Pneumonia, Grad-shap
Case: Pneumonia, Integrated Gradients
Figure 8: For each plot, LDP explanation is on Left and Vanilla explanation is on Right
Case: Pneumonia, SmoothGrad
Case: Pneumonia, Saliency
Case: TB, Grad-shap
Case: TB, Integrated Gradients
Case: TB, SmoothGrad
Case: TB, Saliency
Figure 9: For each plot, LDP explanation is on Left and Vanilla explanation is on Right
Appendix B
We trained ResNet-34 and DenseNet-121 models on the CIFAR-10 dataset using three $\epsilon$ values ($4,7,10$), as lower $\epsilon$ values resulted in harsh privacy-utility trade-off. We weren’t able to train EfficientNet-v2 due to its exuberant computational requirement. We evaluated the models on the whole test set, consisting of total of $10,000$ images. We train all models (both DP and non-DP counterparts) with the exact set of hyperparameters over $50$ epochs. For brevity, we report $Acc_{\mathcal{M}}$ (Table 2), $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}$, and ×$\times$-- × - taking all classes together for Perf Comp. and Agreement in Figure 12.
However, in this case as well Integrated Gradients and Grad-Shap yield $30\%$ as mean DS score and the rest of the explainers also do not obtain $PIS_{Avg}>0.3$ (Figure 12). Furthermore, here also we do not find any (apparent) $PIS_{Avg}$ follows with Perf Comp. and/or Agreement. All in all, the results from our primary experiment are sufficiently comparable here as well. We will release the weights of these models upon publication.
For ResNet-34, we observed that sensitivity is independent across all layers for $\epsilon=10$. However, only the last 2-3 layers for other $\epsilon$ values for ResNet-34 exhibited independent sensitivity. It indicates that independent sensitivity even in the last few layers can potentially make the explanations incomparable across $\epsilon$ values. In contrast, for DenseNet-121, independent sensitivity was consistently observed for all layers across all $\epsilon$. For both types of models, we obtained fewer than 5–7% of layers where we couldn’t reject the null hypothesis for independence of representation. We have reported the dCKA heatmap for DenseNet-121 and ResNet-34 models in figure 11, 10 respectively. Notably, the final layer for both ResNet-34 and DesneNet-121 consistently demonstrated independent sensitivity, rendering Grad-CAM unsuitable for generating comparable explanations. Consequently, we focused on other explainability methods in Figure 12.
@@
LIME Ribeiro et al. (2016) approximates a complex model locally using a simpler interpretable model (e.g., linear regression). For an input $x$, LIME generates perturbed versions of $x$ and calculates the corresponding predictions to fit a local surrogate model. Mathematically, LIME solves the following optimization problem:
@@
SHAP Lundberg and Lee (2017) explains model predictions based on cooperative game theory. For a prediction $f(x)$, SHAP attributes contributions to each feature using Shapley values:
@@
SmoothGrad Smilkov et al. (2017) reduces noise in saliency maps by averaging gradients over multiple noisy samples of the input:
@@
Integrated Gradients Sundararajan et al. (2017) attribute feature importance by integrating the gradients along the path from a baseline input $x^{\prime}$ to the actual input $x$:
@@
Grad-CAM Selvaraju et al. (2019) generates heatmaps for convolutional neural networks by using gradients of the target output with respect to feature maps of a convolutional layer. For a given feature map $A_{k}$, the weights are computed as:
@@
Implementation Details for Section 6 & 7
We employ the off-the-shelf, publicly available implementations of the explainers from Captum library. (https://captum.ai). For (d)CKA, we utilized the publicly available package: Simtorch (https://github.com/ykumards/simtorch) with default (hyper)parameter selection. For Statistical testing with HSIC, we utilised the publicly available package: PyRKHSstats (https://github.com/Black-Swan-ICL/PyRKHSstats) with default (hyper)parameter selection except for the default p-value cutoff of $0.01$. We have used p-value threshold of $0.05$ throughout our experiments. In all our experiments, we considered the whole test set at once as a single batch for dCKA calculation and statistical testing with HSIC.

### v4

@@ 10 Related Work
Privacy-preserving machine learning (PPML) and Explainable AI (XAI) are well-studied research areas, and we’ve outlined the topics under consideration for this study in section 2. For a comprehensive review of PPML and XAI, we direct readers to the excellent surveys by Boulemtafes et al. (Boulemtafes et al., 2020) and Saeed et al. (Saeed and Omlin, 2023), respectively.
Privacy-preserving explainable AI (PPXAI). PPXAI methods are primarily an emerging class of XAI methods specifically made to protect sensitive information in their explanation (desirably with provable guarantees), primarily about the training data. PPXAI methods have only recently begun to emerge, with approaches such as differentially private Locally Linear Maps (LLM) (Harder et al., 2020), differentially private feature-based model explanations (Patel et al., 2022), and differentially private counterfactual explanations (Mochaourab et al., 2021; Yang et al., 2022), etc. Harder et al. in (Harder et al., 2020) proposed a family of simple models to approximate complex models using locally linear maps for each class, which aims to provide DP explanations. Patel et al. (Patel et al., 2022) introduced a perturbation-based algorithm that aims to make the explanation dataset DP, using which they, in a black-box fashion, generate a private explanation for a given point-of-interest. Mochaourab et al. (Mochaourab et al., 2021) made a DP Support Vector Machine and introduce methods for generating robust counterfactuals, Yang et al. (Yang et al., 2022) utilises a DP autoencoder that creates privacy-preserving prototypes by perturbing input data to minimize distance to a counterfactual, while promoting a specific class outcome. For a detailed overview of PPXAI, we direct readers to the comprehensive survey by Nguyen et al. (Nguyen et al., 2024).
Our work is completely orthogonal to PPXAI methods. It is worth noting that our work doesn’t propose or analyse any PPXAI method but explains the interplay between off-the-shelf, well-regarded explainers with DP models. We first establish the desiderata for DP explanations and then mechanistically explore why DP models are particularly challenging to explain using traditional post-hoc explainers. We not only investigate global DP and its inherent limitations for RTE but also explore an alternative route to generate useful private explanations using local DP to generate proxy private explanations.
@@
Recently, researchers have begun investigating the quality of explanations in privacy-preserving environments. Bozorgpanah et al. (Bozorgpanah et al., 2022) generated private datasets (benchmark tabular datasets) using masking methods and compared the Shapley values (Lundberg and Lee, 2017) for test instances given the two models: trained on non-private data and trained on private data. They found minimal effect on moderate protection. Lucieri et al. (Lucieri et al., 2023) investigated the effect of DP training on concept-based explanations on biomedical datasets, and found that DP decreases average Concept Activation Vectors (CAV) accuracy and increases standard deviation; whereas (Saifullah et al., 2024) has reported that differential privacy and federated learning may yield ‘noisy’ feature attribution scores for post-hoc explainers. Berning et al. (Berning et al., 2024) found that k-anonymity degrades the quality of counterfactual explanations on a tabular dataset. Bozorgpanah et al. (Bozorgpanah and Torra, 2024) generated private datasets (benchmark tabular datasets) using masking methods and noise addition and applied TreeSHAP to achieve plausible explanations.
Our work shares some similarities with the aforementioned studies, but it differs significantly in approach. While previous research typically uses quality measures from the XAI literature in an ad-hoc manner within privacy-preserving environments, our systematic investigation is grounded in our own proposed postulate and measures: the Localization Assumption (LA) and Privacy Invariance Score (PIS). While earlier works focus on measuring the degradation in explanation quality within privacy-preserving settings using a limited set of explainers, we demonstrate under a tightly controlled environment that the issue isn’t simply any degradation in explanation quality. Rather, it’s the intrinsic nature of DP models themselves that leads to disobeying LA and producing orthogonal explanations across all commonly used gradient-based explainers, and we are the first ones to mechanistically interpret this very phenomenon. Moreover, many prior studies select explainers and quality measures in a rather ad-hoc fashion. In contrast, we acknowledge our apprehension regarding the suitability of model-agnostic explainers for DP models. Additionally, while Lucieri et al. (Lucieri et al., 2023) and Saifullah et al. (Saifullah et al., 2024) have focused on global DP, the majority of existing research has dealt with local DP.
@@
Analyzing network similarity. Analyzing neural network similarity is crucial for interpreting and improving model behavior. This can be broadly classified into two main types: (i) representational similarity (RS), which measures differences in intermediate layer activations, and (ii) functional similarity, which evaluates discrepancies in model outputs (Klabunde et al., 2024).
@@
To investigate why DP models fail to accommodate common post-hoc explainers, we investigated RS to examine layer-wise changes in activations and their sensitivities. However, in such cases, the similarity measures researchers employ adhere to specific invariances: Permutations, Orthogonal Transformations, Isotropic Scaling, Translations, etc, are to name a few. We have considered the most commonly used set of invariances and chose CKA for measuring RS, but CKA, by its design, comes with a few shortcomings as discussed in section 7. Consequently, we chose dCKA proposed by Cui et al. (Cui et al., 2022) for our analysis. For a broader view on neural network similarity, we refer the readers to the wonderful survey by Klabunde et al. (Klabunde et al., 2024).
@@ 11 Conclusion and Future Studies
We started our research with a simple question of whether we can achieve RTP and RTE together. We kept an eye on the pitfalls of commonly used evaluation metrics for explainability and proposed our desiderata, and found that no commonly used gradient-based explainers are useful for private models. We investigate the activations inside a DP model and how the model is sensitive towards those; we discovered that the intrinsic behaviour of DP models is the key reason behind this. Our mechanistic insights of the private models across different privacy guarantees highlight how DP training alters the internal representations and their sensitivities. It gives a fresh perspective on interpreting DP models and their behaviour from a nuanced angle. Lastly, we make use of LDP to achieve private explanations and conclude our study by outlining the pipeline for the industrial software for our use case that respects both RTP and RTE.
In future studies, we aim to go deeper into the sensitivity landscape of DP models by investigating second-order and higher-order derivatives of the model’s output w.r.t. representations, which may be helpful while investigating a few niche explainers’ behaviour (e.g.: Integrated Hessian (Janizek et al., 2021)), which doesn’t directly leverage immediate first-order gradients. However, such explainers are still not widely used in high-stakes unlike the ones we have considered in this paper. While this work elucidates the interplay between the highly regarded gradient-based explainers and DP models, and explains the degradation of explanations of DP models, examining higher-order derivatives could potentially uncover richer structural patterns in the sensitivity space that aren’t immediately apparent through first-order analysis.
@@ Acknowledgement
We are thankful to Jim Conant for the fruitful discussion on high-dimensional geometry. We are thankful to Yogesh Kumar, the co-author of dCKA (Cui et al., 2022), for a productive discussion. We are also thankful to Dr. Venkata Abhinay Talasila from SRM University AP for his immense cooperation throughout.
Appendix A.
We trained ResNet-34 and DenseNet-121 models on the CIFAR-10 dataset using three $\epsilon$ values ($4,7,10$), as lower $\epsilon$ values resulted in harsh privacy-utility trade-off. We weren’t able to train EfficientNet-v2 due to its exuberant computational requirements. We evaluated the models on the whole test set. We train all models (both DP and non-DP counterparts) with the exact set of hyperparameters over $50$ epochs. For brevity, we report $Acc_{\mathcal{M}}$ (Table 4), $Acc_{\mathcal{M^{\prime}}/\mathcal{M}}$, and ×$\times$-- × - taking all classes together for Perf Comp. and Agreement in Figure 10.
However, in this case as well Integrated Gradients and Grad-Shap yield $30\%$ as mean DS score, and the rest of the explainers also do not obtain $PIS_{Avg}>0.3$ (Figure 10). Furthermore, here also we do not find any (apparent) $PIS_{Avg}$ follows with Perf Comp. and/or Agreement. All in all, the results from our primary experiment are sufficiently comparable here as well. We will release the weights of these models upon publication.
For ResNet-34, we observed that sensitivity is independent across all layers for $\epsilon=10$. However, only the last 2-3 layers for other $\epsilon$ values for ResNet-34 exhibited independent sensitivity. It indicates that independent sensitivity, even in the last few layers, can potentially make the explanations incomparable across $\epsilon$ values. In contrast, for DenseNet-121, independent sensitivity was consistently observed for all layers across all $\epsilon$. For both types of models, we obtained fewer than 5–7% of layers where we couldn’t reject the null hypothesis for independence of representation. We have reported the dCKA heatmap for DenseNet-121 and ResNet-34 models in Figure 9, 8, respectively. Notably, the final layer for both ResNet-34 and DesneNet-121 consistently demonstrated independent sensitivity, rendering Grad-CAM unsuitable for generating comparable explanations. Consequently, we focused on other explainability methods in Figure 10.
@@
LIME [Ribeiro et al., 2016] approximates a complex model locally using a simpler interpretable model (e.g., linear regression). For an input $x$, LIME generates perturbed versions of $x$ and calculates the corresponding predictions to fit a local surrogate model. Mathematically, LIME solves the following optimization problem:
@@
SHAP [Lundberg and Lee, 2017] explains model predictions based on cooperative game theory. For a prediction $f(x)$, SHAP attributes contributions to each feature using Shapley values:
@@
SmoothGrad [Smilkov et al., 2017] reduces noise in saliency maps by averaging gradients over multiple noisy samples of the input:
@@
Integrated Gradients [Sundararajan et al., 2017] attribute feature importance by integrating the gradients along the path from a baseline input $x^{\prime}$ to the actual input $x$:
@@
Grad-CAM [Selvaraju et al., 2019] generates heatmaps for convolutional neural networks by using gradients of the target output with respect to feature maps of a convolutional layer. For a given feature map $A_{k}$, the weights are computed as:
@@
Appendix C.
11.1 Additional Details on Experimental Setup
We train the non-private and private models fixing all the hyperparameters (batch size: 128, lr: 0.001, delta (for DP): 0.001) except for the number of epochs, as private models need more computation to learn due to the heavy regularization DP introduces in the training [Ponomareva et al., 2023]. Following [Ponomareva et al., 2023], we initialised all our models (both non-private and private counterparts) with publicly available pre-trained weights (ImageNet) for better convergence. Furthermore, to make a fair comparison we have fixed the number of all hyperparameters in all the private models with different $\epsilon$. We set the no. of epochs as 50 for all private models; it yielded competitive accuracy across model types. However, within $13-16$ epochs, all non-private models achieved accuracy $>95\%$.
However, while training, we discovered an issue in the vanilla architecture of the aforementioned networks: they utilize batch normalization (BatchNorm). Nevertheless, BatchNorm normalizes a sample based on the statistics of the batch it is in. This means the same sample can get different normalized values depending on the other samples in the batch. For differential privacy, each sample’s privacy needs to be independently preserved. Since BatchNorm depends on other samples, it violates this principle and leaks information about other samples. [Yousefpour et al., 2022] advises replacing BatchNorm layers with privacy-friendly options like Group Normalization, Layer Normalization, Instance Normalization, etc. From the engineering perspective, we have to select one such replacement that scales with sufficiently large datasets without hampering the privacy bounds. Based on previous empirical evidence [Subramani et al., 2021], we replace the BatchNorm layers with GroupNorm layers in all non-private models along with their private counterparts, as GroupNorm does not alter the base architecture drastically, scales well, and adheres to the privacy principle strictly. We utilized the Opacus library for DP-training (https://opacus.ai).
We employ the off-the-shelf, publicly available implementations of the explainers from Captum library. (https://captum.ai).
11.2 Implementation Details for Section 7
For (d)CKA, we utilized the publicly available package: Simtorch (https://github.com/ykumards/simtorch) with default (hyper)parameter selection. For Statistical testing with HSIC, we utilised the publicly available package: PyRKHSstats (https://github.com/Black-Swan-ICL/PyRKHSstats) with default (hyper)parameter selection except for the default p-value cutoff of $0.01$. We have used a p-value threshold of $0.05$ throughout our experiments. In all our experiments, we considered the whole test set at once as a single batch for dCKA calculation and statistical testing with HSIC.